# Modified Dai-Liao Spectral Conjugate Gradient Method with Application to Signal Processing

### v2

@@ 4 Numerical Experiments
In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCG) in [28]. All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Windows 11.
For the numerical experiment, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function:
@@
During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG. We take the initial point $x_{0}=(1,0.8)$. During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealefig1 and LABEL:bealefig2. In Figures LABEL:bealefig1 (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively. Figures LABEL:bealefig1 (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively. The graph between different values of sigma and number of iterations is given in Figure LABEL:bealefig2 (a) and the graph between different values of sigma and execution time is in Figure LABEL:bealefig2 (b).
From the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios. As shown in Table 1, Figure LABEL:bealefig1 and LABEL:bealefig2 (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time. Additionally, From Figure LABEL:bealefig2 (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
@@
$f(\mathbf{x})=\frac{1}{2}\mathbf{x}^{T}A\mathbf{x}+\mathbf{b}^{T}\mathbf{x}$
where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbation. This function has a global minimum at $x=-A^{-1}b$. Note that the gradient of $f$ is given by
@@
For the numerical experiments, we have considered three cases, each characterized by dimensions of1000100010001000, $100$ and $25$. To the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
For the case with dimension $25$, we take the initial point
@@
For all the cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$. Additionally, for the MDDLSCG algorithm, we set $p=0.4$ and $q=0.2$.
During the implementation, the iterations are terminated when the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
The performance of the algorithms MDDLSCG, MSCG, and ScCG for the APQ function is summarized in Table 2 for all three cases. To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure LABEL:APQgraph. From Table 2 and Figure LABEL:APQgraph, it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
Figure LABEL:APQtrack illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations for the case where the problem dimension is $25$.
Furthermore, Figure LABEL:APQprogress depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension. It is observed that MDDLSCG demonstrates superior convergence characteristics compared to MSCG and ScCG.
@@
Table 2: Comparison table of MDDLSCG, MSCG and ScCG for APQ function
5 Application in compressed sensing
Compressed sensing is a well-known signal processing strategy for efficiently acquiring and reconstructing signals with a sparse structure, allowing for memoryless and compact storage of the signals [30]. The importance of compressed sensing in real-world applications, such as machine learning, compressive imaging, radar, wireless sensor networks, medical imaging, astrophysical signals, and video coding, has been highlighted in [31].
The compressed sensing problem involves finding a sparse solution to a highly underdetermined system, represented as $Ax=b$, where $A\in\mathbb{R}^{m\times n}$ (with $m\ll n$) and $b\in\mathbb{R}^{m}$ [30]. A common approach is to solve the following optimization problem:
@@ 5 Application in compressed sensing
where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that controls the trade-off between sparsity and reconstruction fidelity. A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [32] and the references therein. However, solving BPD is challenging due to the non-smoothness of the $\ell_{1}$-penalty term. To address this issue, Zhu et al. [33] proposed a relaxation of BPD using Nesterov’s smoothing technique [34], resulting in the following formulation:
@@
where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined as:
@@
$\psi_{\lambda}(|x_{i}|)=\begin{cases}\frac{|x_{i}|^{2}}{2\lambda},&\text{if }|x_{i}|<\lambda,\\ |x_{i}|-\frac{\lambda}{2},&\text{if }|x_{i}|\geq\lambda.\end{cases}$
Here, $f_{\lambda}$ is known as the Huber function [35], and its gradient $\nabla f_{\lambda}$ has a Lipschitz constant of $\frac{1}{\lambda}$ [36], with the explicit form given by:
$(\nabla f_{\lambda})_{i}=\begin{cases}\frac{x_{i}}{\lambda},&\text{if }|x_{i}|<\lambda,\\ \text{sgn}(x_{i}),&\text{otherwise},\end{cases}$
where $\lambda>0$ is a constant and $\text{sgn}(\cdot)$ denotes the signum function.
Therefore, instead of solving the original problem $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\|x\|_{1}$, we solve the smoothed version:
@@
[…]
@@
To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, which is defined as follows:
@@
Let $k$ denote the number of nonzero elements in the original signal. For a fair comparison, we consider two different settings: $(m,n,k)=(128,512,16)$, $(256,1024,32)$.
The detailed numerical results are presented in Table 3, where we report the average number of iterations (Itr), average CPU time (Tcpu), mean squared error (MSE), and relative error (RelErr) for each setting. The graph depicting the relationship between relative error and the number of iterations is shown in Figures LABEL:stpp512(a) and LABEL:stopp(a), while the graph illustrating MSE versus the number of iterations is presented in Figures LABEL:stpp512(b) and LABEL:stopp(b). Additionally, the recovery of the signal is visualized in Figures 11 and 13.
From Table 3, Figures LABEL:stpp512, and LABEL:stopp, it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
@@
Dim$(m,n,k)$ Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
@@ 6 Conclusions
We investigated a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction. The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques. Theoretical analysis establishes the global convergence of the method for general nonlinear functions under standard assumptions. Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm. Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [26] and the ScCG method [28], highlighting its potential in optimization and related applications.
@@
Acknowledgments
The authors are grateful to the editor and to the reviewers for their remarks and advice which improved our manuscript.

### v3

@@ 4 Numerical Experiments
In this section we present the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [Faramarzi2019] (denoted by MSCG) and the scaled conjugate gradient method (denoted by ScCG) in [Mrad2024]. All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Windows 11.
For the numerical experiments, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function.
@@
(a) Surface and Contour plot of the Beale functions
For the numerical tests, we choose $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG. We take the initial point $x_{0}=(1,0.8)$. The iterations are terminated once they satisfy the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure 8(a) and 8(b). In Figures 8(a) (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively. Figures 8(a) (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively. The comparison graph between different values of sigma and the number of iterations is depicted in Figure 8(b) (a) and the comparison graph between different values of sigma and execution time is in Figure 8(b) (b).
In view of the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios. As shown in Table 1, Figure 8(a) and 8(b) (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time. Additionally, using Figure 8(b) (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
@@
(a) Comparison of the iterations and execution time for different $\sigma$ for MDDLSCG, MSCG, ScCG for the Beale function
(b) Comparison of the iterations and execution time for different $\sigma$ for MDDLSCG, MSCG, ScCG for the Beale function
@@
$f(x)=\frac{1}{2}x^{T}Ax+b^{T}x\text{ for all }x\in\mathbb{R}^{m},$
where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbations. This function has a global minimum at $x=-A^{-1}b$. Note that the gradient of $f$ is given by
@@
For the numerical experiments, we have considered three cases, each characterized by dimensions of $1000$, 100100 and $25$. In the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
In the case of dimension $25$, we take the initial point
@@
In all cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$. Additionally, for the MDDLSCG algorithm, we set $p=0.4$ and $q=0.2$.
The iterations are terminated once they satisfy the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
The performance of the algorithms MDDLSCG, MSCG, and ScCG for the APQ function is summarized in Table 2 for all three cases. To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure 11(a). From Table 2 and Figure 11(a), it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
Figure 11(b) illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations in the case where the problem dimension is $25$.
Furthermore, Figure 11(c) depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension. It is observed that MDDLSCG demonstrates superior convergence characteristics compared to MSCG and ScCG.
@@
Table 2: Comparison table of MDDLSCG, MSCG and ScCG for the APQ function
(a) Convergence behavior of $\|\nabla f(x_{n})\|_{\infty}$ for MDDLSCG, MSCG, ScCG for the APQ function
(b) Tracking behaviour of MDDLSCG, MSCG, ScCG for the APQ function at various iterations
(c) Graphical progression of MDDLSCG, MSCG, ScCG upto 50th iterations for the APQ function
5 Application to compressed sensing
Compressed sensing is a widely recognized signal processing approach that enables the efficient acquisition and reconstruction of signals with a sparse structure, allowing these signals to be stored in a compact and memoryless form [Bruckstein2009]. Its practical significance has been demonstrated in numerous fields, including machine learning, compressive imaging, radar systems, wireless sensor networks, medical and astronomical signal processing, and video compression, as discussed in [Aminifard2022].
The compressed sensing problem involves finding a sparse solution to a highly underdetermined system, represented as $Ax=b$, where $A\in\mathbb{R}^{m\times n}$ (with $m\ll n$) and $b\in\mathbb{R}^{m}$ [Bruckstein2009]. A common approach is to solve the following optimization problem:
@@ 5 Application to compressed sensing
where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that regulates the trade-off between sparsity and reconstruction fidelity. A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [Esmaeili2019] and references therein. However, solving BPD is difficult because the $\ell_{1}$-penalty term is non-smooth. To overcome this difficulty, Zhu et al. [Zhu2013] introduced a relaxed form of BPD based on Nesterov’s smoothing technique [Nesterov2005Siam], leading to the following formulation:
@@
where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined by
@@
$\psi_{\lambda}(|x_{i}|)=\begin{cases}\frac{|x_{i}|^{2}}{2\lambda},&\text{if }|x_{i}|<\lambda,\\ |x_{i}|-\frac{\lambda}{2},&\text{if }|x_{i}|\geq\lambda,\end{cases}$
where $\lambda>0$. Function $f_{\lambda}$ is known as the Huber function [Rangarajan], and its gradient $\nabla f_{\lambda}$ is $\frac{1}{\lambda}$ -Lipschitz continuous [Becker2011]. The explicit form of the Huber function is given by
[…]
where $\text{sgn}(\cdot)$ denotes the signum function.
Therefore, instead of solving the original problem for $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\|x\|_{1}$, we solve the smoothed version:
@@
For MDDLSCG, we choose $p=0.4$ and $q=0.2$. The Gaussian matrix $A$ is generated with the MATLAB commands randn(m, n) and the noise vector $w$ is generated with $0.1\times\texttt{randn(m, 1)}$. The measurement vector $b$ is computed by $b=Ax+w.$ The regularization parameter is chosen as $\mu=0.001\|A^{T}b\|_{\infty}$. Each methods begin with the initial point $x_{0}=A^{T}b$. Let $x$ represent the true signal and $x^{*}$ denote the recovered one. The termination criterion for the experiment is the mean-squared error(MSE):
@@
To assess the reconstruction quality, we also compute the relative error (RelErr) of the restored signal, defined by
@@
Let $k$ represent the number of nonzero components in the original signal. To ensure a fair comparison, we test two cases: $(m,n,k)=(128,512,16)$, $(256,1024,32)$.
The numerical results are summarized in Table 3, where we list, for each case, the average iteration count (Itr), the average CPU time (Tcpu), the mean squared error (MSE), and the relative error (RelErr). The graph depicting the relationship between relative error and the number of iterations is shown in Figures 11(d)(a) and 13(a)(a), while the graph illustrating MSE versus the number of iterations is presented in Figures 11(d)(b) and 13(a)(b). In addition, the recovery of the signal is visualized in Figures 11 and 13.
From Table 3, Figures 11(d) and 13(a), it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
(d) Convergence behavior and relative error plots, demonstrating the performance of the methods MDDLSCG, MSCG and ScCG for $(m,n,k)=(128,512,16)$
@@
(a) Convergence behavior and relative error plots, demonstrating the performance of the methods MDDLSCG, MSCG and ScCG for $(m,n,k)=(256,1024,32)$
@@
Dim(m,n,k)(m,n,k) Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
@@ 6 Conclusions
We investigated a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction. The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques. Theoretical analysis establishes the global convergence of our method for general nonlinear functions under standard assumptions. Numerical experiments and an application to signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm. Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [Faramarzi2019] and the ScCG method [Mrad2024], highlighting its potential in optimization and related applications.